# Key Words.: Galaxies: clusters: general – Galaxies: clusters: individual: Abell 426 – X-rays: galaxies: clusters – Radio continuum: galaxies – Galaxies: clusters: intracluster medium – Galaxies: kinematics and dynamics1 Introduction

### v1

@@ 4 Conclusions
At larger scales, the excess X-ray emission can be traced westwards up to the virial radius, with some tentative arc-like features in the NW direction. It is plausible that these structures are associated with the same merger.
Finally, we discussed briefly the case of the high-velocity radio galaxy NGC1265, known for its curious radio tail, which also shows sharp bends. In our ”minimalist” model, this galaxy is moving nearly along the line of sight with the total 3D velocity and is currently on the far side of the main cluster. The gas motions induced by IC310 merger might play a role in shaping its tail, although other explanations are possible, too.
@@ Acknowledgements.
This work is partly based on observations with the eROSITA telescope onboard SRG space observatory. The SRG observatory was built by Roskosmos in the interests of the Russian Academy of Sciences, represented by its Space Research Institute (IKI), in the framework of the Russian Federal Space Program, with the participation of the Deutsches Zentrum für Luft- und Raumfahrt (DLR). The eROSITA X-ray telescope was built by a consortium of German Institutes led by MPE and supported by DLR. The SRG spacecraft was designed, built, launched, and operated by the Lavochkin Association and its subcontractors. The science data are downlinked via the Deep Space Network Antennae in Bear Lakes, Ussurijsk, and Baikonur, funded by Roskosmos. The development and construction of the eROSITA X-ray instrument was led by MPE, with contributions from the Dr. Karl Remeis Observatory Bamberg ECAP (FAU Erlangen-Nuernberg), the University of Hamburg Observatory, the Leibniz Institute for Astrophysics Potsdam (AIP), and the Institute for Astronomy and Astrophysics of the University of Tübingen, with the support of DLR and the Max Planck Society. The Argelander Institute for Astronomy of the University of Bonn and the Ludwig Maximilians Universität Munich also participated in the science preparation for eROSITA. The eROSITA data were processed using the eSASS/NRTA software system developed by the German eROSITA consortium and analyzed using proprietary data reduction software developed by the Russian eROSITA Consortium. This publication makes use of data products from the Two Micron All Sky Survey, which is a joint project of the University of Massachusetts and the Infrared Processing and Analysis Center/California Institute of Technology, funded by the National Aeronautics and Space Administration and the National Science Foundation. IK and KD acknowledges support by the COMPLEX project from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program grant agreement ERC-2019-AdG 882679. The calculations for the hydrodynamical simulations were carried out at the Leibniz Supercomputer Center (LRZ) under the project pr83li (Magneticum).References
- Aleksić et al. (2014) Aleksić, J., Ansoldi, S., Antonelli, L. A., et al. 2014, Science, 346, 1080
- Andreon (1994) Andreon, S. 1994, A&A, 284, 801
- Aung et al. (2023) Aung, H., Nagai, D., Rozo, E., Wolfe, B., & Adhikari, S. 2023, MNRAS, 521, 3981
- Beck et al. (2016) Beck, A. M., Murante, G., Arth, A., et al. 2016, MNRAS, 455, 2110
- Begelman et al. (1979) Begelman, M. C., Rees, M. J., & Blandford, R. D. 1979, Nature, 279, 770
- Bellomi et al. (2024) Bellomi, E., ZuHone, J. A., Weinberger, R., et al. 2024, ApJ, 974, 234
- Biffi et al. (2022) Biffi, V., Dolag, K., Reiprich, T. H., et al. 2022, A&A, 661, A17
- Birnboim et al. (2010) Birnboim, Y., Keshet, U., & Hernquist, L. 2010, MNRAS, 408, 199
- Boehringer et al. (1993) Boehringer, H., Voges, W., Fabian, A. C., Edge, A. C., & Neumann, D. M. 1993, MNRAS, 264, L25
- Branduardi-Raymont et al. (1981) Branduardi-Raymont, G., Fabricant, D., Feigelson, E., et al. 1981, ApJ, 248, 55
- Brentjens (2011) Brentjens, M. A. 2011, A&A, 526, A9
@@
- Brunner et al. (2018) Brunner, H., Boller, T., Coutinho, D., et al. 2018, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 10699, Space Telescopes and Instrumentation 2018: Ultraviolet to Gamma Ray, ed. J.-W. A. den Herder, S. Nikzad, & K. Nakazawa, 106995G
- Carucci et al. (2014) Carucci, I. P., Sparre, M., Hansen, S. H., & Joyce, M. 2014, J. Cosmology Astropart. Phys, 2014, 057
- Chandran et al. (2023) Chandran, J., Remazeilles, M., & Barreiro, R. B. 2023, MNRAS, 526, 5682
- Churazov et al. (2000) Churazov, E., Forman, W., Jones, C., & Böhringer, H. 2000, A&A, 356, 788
- Churazov et al. (2003) Churazov, E., Forman, W., Jones, C., & Böhringer, H. 2003, ApJ, 590, 225
- Churazov et al. (2023) Churazov, E., Khabibullin, I., Bykov, A. M., Lyskova, N., & Sunyaev, R. 2023, A&A, 670, A156
- Churazov et al. (2021) Churazov, E., Khabibullin, I., Lyskova, N., Sunyaev, R., & Bykov, A. M. 2021, A&A, 651, A41
- de Bruyn & Brentjens (2005) de Bruyn, A. G. & Brentjens, M. A. 2005, A&A, 441, 931
- Dolag et al. (2016) Dolag, K., Komatsu, E., & Sunyaev, R. 2016, MNRAS, 463, 1797
- Dolag et al. (2025) Dolag, K., Remus, R.-S., Valenzuela, L. M., et al. 2025, arXiv e-prints, arXiv:2504.01061
- Dunn et al. (2010) Dunn, R. J. H., Allen, S. W., Taylor, G. B., et al. 2010, MNRAS, 404, 180
- Ettori et al. (1998) Ettori, S., Fabian, A. C., & White, D. A. 1998, MNRAS, 300, 837
- Fabian et al. (2011) Fabian, A. C., Sanders, J. S., Allen, S. W., et al. 2011, MNRAS, 418, 2154
- Fabian et al. (2003) Fabian, A. C., Sanders, J. S., Allen, S. W., et al. 2003, MNRAS, 344, L43
- Gendron-Marsolais et al. (2020) Gendron-Marsolais, M., Hlavacek-Larrondo, J., van Weeren, R. J., et al. 2020, MNRAS, 499, 5791
- Graham et al. (2019) Graham, J. A., Brown, A. M., & Chadwick, P. M. 2019, MNRAS, 485, 3277
- Green et al. (2019) Green, G. M., Schlafly, E., Zucker, C., Speagle, J. S., & Finkbeiner, D. 2019, ApJ, 887, 93
- Groth et al. (2023) Groth, F., Steinwandel, U. P., Valentini, M., & Dolag, K. 2023, MNRAS, 526, 616
- Güver & Özel (2009) Güver, T. & Özel, F. 2009, MNRAS, 400, 2050
- Hirschmann et al. (2014) Hirschmann, M., Dolag, K., Saro, A., et al. 2014, MNRAS, 442, 2304
- Hitomi Collaboration et al. (2016) Hitomi Collaboration, Aharonian, F., Akamatsu, H., et al. 2016, Nature, 535, 117
- Huchra et al. (2012) Huchra, J. P., Macri, L. M., Masters, K. L., et al. 2012, ApJS, 199, 26
- HyeongHan et al. (2025) HyeongHan, K., Jee, M. J., Lee, W., et al. 2025, Nature Astronomy, 9, 925
- Kadler et al. (2012) Kadler, M., Eisenacher, D., Ros, E., et al. 2012, A&A, 538, L1
- Kang et al. (2024) Kang, W., Hwang, H. S., Song, H., et al. 2024, ApJS, 272, 22
- Kent & Sargent (1983) Kent, S. M. & Sargent, W. L. W. 1983, AJ, 88, 697
- Khabibullin et al. (2023) Khabibullin, I. I., Churazov, E. M., Bykov, A. M., Chugai, N. N., & Sunyaev, R. A. 2023, MNRAS, 521, 5536
- Kluge et al. (2025) Kluge, M., Hatch, N. A., Montes, M., et al. 2025, A&A, 697, A13
- Komatsu et al. (2011) Komatsu, E., Smith, K. M., Dunkley, J., et al. 2011, ApJS, 192, 18
- Lal et al. (2022) Lal, D. V., Lyskova, N., Zhang, C., et al. 2022, ApJ, 934, 170
- Lyskova et al. (2019) Lyskova, N., Churazov, E., Zhang, C., et al. 2019, MNRAS, 485, 2922
- Markevitch & Vikhlinin (2007) Markevitch, M. & Vikhlinin, A. 2007, Phys. Rep., 443, 1
- Marleau et al. (2025) Marleau, F. R., Cuillandre, J. C., Cantiello, M., et al. 2025, A&A, 697, A12
- Meusinger et al. (2020) Meusinger, H., Rudolf, C., Stecklum, B., et al. 2020, A&A, 640, A30
- Miley et al. (1975) Miley, G. K., Wellington, K. J., & van der Laan, H. 1975, A&A, 38, 381
- Navarro et al. (1997) Navarro, J. F., Frenk, C. S., & White, S. D. M. 1997, ApJ, 490, 493
- Pfrommer & Jones (2011) Pfrommer, C. & Jones, T. W. 2011, ApJ, 730, 22
- Predehl et al. (2021) Predehl, P., Andritschke, R., Arefiev, V., et al. 2021, A&A, 647, A1
- Predehl & Schmitt (1995) Predehl, P. & Schmitt, J. H. M. M. 1995, A&A, 293, 889
- Roberts & Parker (2020) Roberts, I. D. & Parker, L. C. 2020, MNRAS, 495, 554
- Roediger et al. (2024) Roediger, E., Vaezzadeh, I., & Nulsen, P. 2024, MNRAS, 529, 563
- Rudnick (2021) Rudnick, L. 2021, Galaxies, 9, 85
- Ryle & Windram (1968) Ryle, M. & Windram, M. D. 1968, MNRAS, 138, 1
- Sales et al. (2007) Sales, L. V., Navarro, J. F., Abadi, M. G., & Steinmetz, M. 2007, MNRAS, 379, 1475
- Sanders et al. (2020) Sanders, J. S., Dennerl, K., Russell, H. R., et al. 2020, A&A, 633, A42
- Sato et al. (2005) Sato, K., Furusho, T., Yamasaki, N. Y., et al. 2005, PASJ, 57, 743
- Schwarz et al. (1992) Schwarz, R. A., Edge, A. C., Voges, W., et al. 1992, A&A, 256, L11
- Sheardown et al. (2019) Sheardown, A., Fish, T. M., Roediger, E., et al. 2019, ApJ, 874, 112
- Sijbring & de Bruyn (1998) Sijbring, D. & de Bruyn, A. G. 1998, A&A, 331, 901
- Simionescu et al. (2011) Simionescu, A., Allen, S. W., Mantz, A., et al. 2011, Science, 331, 1576
- Simionescu et al. (2012) Simionescu, A., Werner, N., Urban, O., et al. 2012, ApJ, 757, 182
- Springel (2005) Springel, V. 2005, MNRAS, 364, 1105
- Sunyaev et al. (2021) Sunyaev, R., Arefiev, V., Babyshkin, V., et al. 2021, A&A, 656, A132
- Tamura et al. (2014) Tamura, T., Yamasaki, N. Y., Iizuka, R., et al. 2014, ApJ, 782, 38
- van Weeren et al. (2024) van Weeren, R. J., Timmerman, R., Vaidya, V., et al. 2024, A&A, 692, A12
- Vikhlinin et al. (2005) Vikhlinin, A., Markevitch, M., Murray, S. S., et al. 2005, ApJ, 628, 655
- Walker et al. (2022) Walker, S. A., Mirakhor, M. S., ZuHone, J., et al. 2022, ApJ, 929, 37
- Zeldovich & Sunyaev (1969) Zeldovich, Y. B. & Sunyaev, R. A. 1969, Ap&SS, 4, 301
- Zhang et al. (2020) Zhang, C., Churazov, E., Dolag, K., Forman, W. R., & Zhuravleva, I. 2020, MNRAS, 498, L130
- Zhang et al. (2019) Zhang, C., Churazov, E., Forman, W. R., & Lyskova, N. 2019, MNRAS, 488, 5259
- Zhu et al. (2021) Zhu, Z., Simionescu, A., Akamatsu, H., et al. 2021, A&A, 652, A147
- Zhuravleva et al. (2015) Zhuravleva, I., Churazov, E., Arévalo, P., et al. 2015, MNRAS, 450, 4184
- Zuhone & Roediger (2016) Zuhone, J. A. & Roediger, E. 2016, Journal of Plasma Physics, 82, 535820301
Appendix A Chandra and XMM datasets

### v2

@@ 4 Conclusions
At larger scales, the excess X-ray emission can be traced westwards up to the virial radius, with some tentative arc-like features in the NW direction. It is plausible that these structures are associated with the same merger. Independent of the details, the elongation of the Perseus cluster X-ray emission in the East-West direction coincides well with the orientation of the large-scale filament of the Perseus-Pisces supercluster.
While this paper was under review, two new studies of the Perseus cluster based on the XRISM observations (XRISM Collaboration et al., 2025; Zhang et al., 2025) became available. Broadly, these studies are consistent with the scenario discussed here.
Finally, we discussed briefly the case of the high-velocity radio galaxy NGC1265, known for its curious radio tail, which also shows sharp bends. In our ”minimalist” model, this galaxy is moving nearly along the line of sight with the total 3D velocity and is currently on the far side of the main cluster. The gas motions induced by the IC310 merger might play a role in shaping its tail, although other explanations are possible, too.
@@ Acknowledgements.
This work is partly based on observations with the eROSITA telescope onboard SRG space observatory. The SRG observatory was built by Roskosmos in the interests of the Russian Academy of Sciences, represented by its Space Research Institute (IKI), in the framework of the Russian Federal Space Program, with the participation of the Deutsches Zentrum für Luft- und Raumfahrt (DLR). The eROSITA X-ray telescope was built by a consortium of German Institutes led by MPE and supported by DLR. The SRG spacecraft was designed, built, launched, and operated by the Lavochkin Association and its subcontractors. The science data are downlinked via the Deep Space Network Antennae in Bear Lakes, Ussurijsk, and Baikonur, funded by Roskosmos. The development and construction of the eROSITA X-ray instrument was led by MPE, with contributions from the Dr. Karl Remeis Observatory Bamberg ECAP (FAU Erlangen-Nuernberg), the University of Hamburg Observatory, the Leibniz Institute for Astrophysics Potsdam (AIP), and the Institute for Astronomy and Astrophysics of the University of Tübingen, with the support of DLR and the Max Planck Society. The Argelander Institute for Astronomy of the University of Bonn and the Ludwig Maximilians Universität München also participated in the science preparation for eROSITA. The eROSITA data were processed using the eSASS/NRTA software system developed by the German eROSITA consortium and analyzed using proprietary data reduction software developed by the Russian eROSITA Consortium. This publication makes use of data products from the Two Micron All Sky Survey, which is a joint project of the University of Massachusetts and the Infrared Processing and Analysis Center/California Institute of Technology, funded by the National Aeronautics and Space Administration and the National Science Foundation. IK and KD acknowledge support by the COMPLEX project from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program grant agreement ERC-2019-AdG 882679. The calculations for the hydrodynamical simulations were carried out at the Leibniz Supercomputer Center (LRZ) under the project pr83li (Magneticum).References
- Black hole lightning due to particle acceleration at subhorizon scales. Science 346 (6213), pp. 1080–1084. External Links: Document, 1412.4936, ADS entry Cited by: §3.3.1.
- Morphological segregation of galaxies in clusters I. The non-uniform distribution of types in the inner region of Perseus.. A&A 284, pp. 801–809. External Links: ADS entry Cited by: §3.1.
- Accurate model of the projected velocity distribution of galaxies in dark matter haloes. MNRAS 521 (3), pp. 3981–3990. External Links: Document, 2204.13131, ADS entry Cited by: §3.3.4.
- An improved SPH scheme for cosmological simulations. MNRAS 455 (2), pp. 2110–2130. External Links: Document, 1502.07358, ADS entry Cited by: §3.3.4.
- A twin-jet model for radio trails. Nature 279, pp. 770–773. External Links: Document, ADS entry Cited by: §3.3.1.
- On the Origin of the Ancient, Large-scale Cold Front in the Perseus Cluster of Galaxies. ApJ 974 (2), pp. 234. External Links: Document, 2310.09422, ADS entry Cited by: §3.1, §3.1.
- The eROSITA view of the Abell 3391/95 field: Case study from the Magneticum cosmological simulation. A&A 661, pp. A17. External Links: Document, 2106.14542, ADS entry Cited by: §3.3.4.
- Cold fronts by merging of shocks. MNRAS 408 (1), pp. 199–212. External Links: Document, 1006.1892, ADS entry Cited by: §3.2.
- A ROSAT HRI study of the interaction of the X-ray emitting gas and radio lobes of NGC 1275.. MNRAS 264, pp. L25–L28. External Links: Document, ADS entry Cited by: §1.
- The Cosmic Large-Scale Structure in X-rays (CLASSIX) Cluster Survey. III. The Perseus-Pisces supercluster and the Southern Great Wall as traced by X-ray luminous galaxy clusters. A&A 651, pp. A16. External Links: Document, 2105.14051, ADS entry Cited by: Figure 1, §2.1.
- Soft X-ray images of the central region of the Perseus cluster.. ApJ 248, pp. 55–60. External Links: Document, ADS entry Cited by: §1, §2.4, §3.1.
- Wide field polarimetry around the Perseus cluster at 350 MHz. A&A 526, pp. A9. External Links: Document, 1011.0854, ADS entry Cited by: §3.3.3.
@@
- eROSITA ground operations. In Space Telescopes and Instrumentation 2018: Ultraviolet to Gamma Ray, J. A. den Herder, S. Nikzad, and K. Nakazawa (Eds.), Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 10699, pp. 106995G. External Links: Document, ADS entry Cited by: §2.
- Particle ejection during mergers of dark matter halos. J. Cosmology Astropart. Phys 2014 (6), pp. 057. External Links: Document, 1405.6725, ADS entry Cited by: §3.3.3.
- An improved Compton parameter map of thermal Sunyaev-Zeldovich effect from Planck PR4 data. MNRAS 526 (4), pp. 5682–5698. External Links: Document, 2305.10193, ADS entry Cited by: Figure 1.
- Asymmetric, arc minute scale structures around NGC 1275. A&A 356, pp. 788–794. External Links: Document, astro-ph/0002375, ADS entry Cited by: §1.
- XMM-Newton Observations of the Perseus Cluster. I. The Temperature and Surface Brightness Structure. ApJ 590 (1), pp. 225–237. External Links: Document, astro-ph/0301482, ADS entry Cited by: Appendix A, §1, §2.4, §3.1, §3.1.
- Tempestuous life beyond R500: X-ray view on the Coma cluster with SRG/eROSITA. II. Shock and relic. A&A 670, pp. A156. External Links: Document, 2205.07511, ADS entry Cited by: Appendix A, Figure 14.
- Tempestuous life beyond R500: X-ray view on the Coma cluster with SRG/eROSITA. I. X-ray morphology, recent merger, and radio halo connection. A&A 651, pp. A41. External Links: Document, 2012.11627, ADS entry Cited by: Appendix A, §2.
- SRG/eROSITA discovery of a large circular SNR candidate G116.6-26.1: SN Ia explosion probing the gas of the Milky Way halo?. MNRAS 507 (1), pp. 971–982. External Links: Document, 2106.09454, ADS entry Cited by: Appendix A.
- Diffuse polarized emission associated with the Perseus cluster. A&A 441 (3), pp. 931–947. External Links: Document, astro-ph/0507351, ADS entry Cited by: §3.3.3.
- SZ effects in the Magneticum Pathfinder simulation: comparison with the Planck, SPT, and ACT results. MNRAS 463 (2), pp. 1797–1811. External Links: Document, 1509.05134, ADS entry Cited by: §3.3.4.
- Encyclopedia Magneticum: Scaling Relations from Cosmic Dawn to Present Day. arXiv e-prints, pp. arXiv:2504.01061 (A&A, submitted). External Links: Document, 2504.01061, ADS entry Cited by: §3.3.4.
- The radio properties of a complete, X-ray selected sample of nearby, massive elliptical galaxies. MNRAS 404 (1), pp. 180–197. External Links: Document, 1001.1630, ADS entry Cited by: §3.1.
- ROSATPSPC observations of the outer regions of the Perseus cluster of galaxies. MNRAS 300 (3), pp. 837–856. External Links: Document, astro-ph/9806375, ADS entry Cited by: §1, §2.4.
- A wide Chandra view of the core of the Perseus cluster. MNRAS 418 (4), pp. 2154–2164. External Links: Document, 1105.5025, ADS entry Cited by: §1, §2.4, §3.1.
- A deep Chandra observation of the Perseus cluster: shocks and ripples. MNRAS 344 (3), pp. L43–L47. External Links: Document, astro-ph/0306036, ADS entry Cited by: §1.
- SRG/eROSITA in-flight background at L2. In Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, J. A. den Herder, S. Nikzad, and K. Nakazawa (Eds.), Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 11444, pp. 114441O. External Links: Document, ADS entry Cited by: §2.
- High-resolution VLA low radio frequency observations of the Perseus cluster: radio lobes, mini-halo, and bent-jet radio galaxies. MNRAS 499 (4), pp. 5791–5805. External Links: Document, 2005.12298, ADS entry Cited by: Figure 7, §3.3.1, §3.3.2, §3.3.
- Fermi-LAT observations of extreme spectral variability in IC 310. MNRAS 485 (3), pp. 3277–3287. External Links: Document, 1903.07897, ADS entry Cited by: §3.1.
- A 3D Dust Map Based on Gaia, Pan-STARRS 1, and 2MASS. ApJ 887 (1), pp. 93. External Links: Document, 1905.02734, ADS entry Cited by: Appendix D, Figure 1, §2.1.
- The cosmological simulation code OPENGADGET3 - implementation of meshless finite mass. MNRAS 526 (1), pp. 616–644. External Links: Document, 2301.03612, ADS entry Cited by: §3.3.4.
- The relation between optical extinction and hydrogen column density in the Galaxy. MNRAS 400 (4), pp. 2050–2053. External Links: Document, 0903.2057, ADS entry Cited by: §2.1.
- Cosmological simulations of black hole growth: AGN luminosities and downsizing. MNRAS 442 (3), pp. 2304–2324. External Links: Document, 1308.0333, ADS entry Cited by: §3.3.4.
- The quiescent intracluster medium in the core of the Perseus cluster. Nature 535 (7610), pp. 117–121. External Links: Document, 1607.04487, ADS entry Cited by: §1.
- The 2MASS Redshift Survey—Description and Data Release. ApJS 199 (2), pp. 26. External Links: Document, 1108.0669, ADS entry Cited by: §2.2.
- Direct evidence of a major merger in the Perseus cluster. Nature Astronomy 9, pp. 925–931. External Links: Document, 2405.00115, ADS entry Cited by: Figure 6, §3.1, §3.1, footnote 3.
- The blazar-like radio structure of the TeV source IC 310. A&A 538, pp. L1. External Links: Document, 1201.2870, ADS entry Cited by: §3.3.1, footnote 6.
- A Deep Redshift Survey of the Perseus Cluster (A426): Spatial Distribution and Kinematics of Galaxies. ApJS 272 (1), pp. 22. External Links: Document, 2403.19307, ADS entry Cited by: §3.1.
- The dynamics of rich clusters of galaxies. II. The Perseus cluster.. AJ 88, pp. 697–708. External Links: Document, ADS entry Cited by: §3.1.
- SRG/eROSITA discovery of a radio-faint X-ray candidate supernova remnant SRGe J003602.3+605421 = G121.1-1.9. MNRAS 521 (4), pp. 5536–5556. External Links: Document, 2207.00064, ADS entry Cited by: Appendix A, §2.
- SRG/eROSITA view of X-ray reflection in the Central Molecular Zone: a snapshot in September-October 2019. MNRAS 509 (4), pp. 6068–6076. External Links: Document, 2111.07899, ADS entry Cited by: Appendix A.
- Study of X-ray emission from the S147 nebula by SRG/eROSITA: Supernova-in-the-cavity scenario. A&A 689, pp. A278. External Links: Document, 2401.17261, ADS entry Cited by: Appendix A.
- Euclid: Early Release Observations – The intracluster light and intracluster globular clusters of the Perseus cluster. A&A 697, pp. A13. External Links: Document, 2405.13503, ADS entry Cited by: §3.1.
- Seven-year Wilkinson Microwave Anisotropy Probe (WMAP) Observations: Cosmological Interpretation. ApJS 192 (2), pp. 18. External Links: Document, 1001.4538, ADS entry Cited by: §3.3.4.
- High-resolution, High-sensitivity, Low-frequency uGMRT View of Coma Cluster of Galaxies. ApJ 934 (2), pp. 170. External Links: Document, 2207.06624, ADS entry Cited by: §3.3.2.
- Close-up view of an ongoing merger between the NGC 4839 group and the Coma cluster - a post-merger scenario. MNRAS 485 (2), pp. 2922–2934. External Links: Document, 1811.07944, ADS entry Cited by: §3.3.1, footnote 2.
- Shocks and cold fronts in galaxy clusters. Phys. Rep. 443 (1), pp. 1–53. External Links: Document, astro-ph/0701821, ADS entry Cited by: §1.
- Euclid: Early Release Observations – Dwarf galaxies in the Perseus galaxy cluster. A&A 697, pp. A12. External Links: Document, 2405.13502, ADS entry Cited by: §3.1.
- The galaxy population within the virial radius of the Perseus cluster. A&A 640, pp. A30. External Links: Document, 2006.07631, ADS entry Cited by: footnote 1.
- The structure of the radio galaxy NGC 1265.. A&A 38 (3), pp. 381–390. External Links: ADS entry Cited by: §3.3.2.
- A Universal Density Profile from Hierarchical Clustering. ApJ 490 (2), pp. 493–508. External Links: Document, astro-ph/9611107, ADS entry Cited by: §3.3.3.
- Radio Galaxy NGC 1265 Unveils the Accretion Shock Onto the Perseus Galaxy Cluster. ApJ 730 (1), pp. 22. External Links: Document, 1004.3540, ADS entry Cited by: §3.3.2, §3.3.3.
- The eROSITA X-ray telescope on SRG. A&A 647, pp. A1. External Links: Document, 2010.03477, ADS entry Cited by: §1, §2.
- X-raying the interstellar medium: ROSAT observations of dust scattering halos.. A&A 293, pp. 889–905. External Links: ADS entry Cited by: §2.1.
- Ram pressure stripping candidates in the coma cluster: evidence for enhanced star formation. MNRAS 495 (1), pp. 554–569. External Links: Document, 2004.12033, ADS entry Cited by: §3.3.2.
- A toy model for gas sloshing in galaxy clusters. MNRAS 529 (1), pp. 563–574. External Links: Document, 2402.18632, ADS entry Cited by: §3.1.
- Radio Galaxy Classification: #Tags, Not Boxes. Galaxies 9 (4), pp. 85. External Links: Document, 2110.13733, ADS entry Cited by: §3.3.2.
- The radio emission from galaxies in the Perseus cluster. MNRAS 138, pp. 1. External Links: Document, ADS entry Cited by: §3.3.2.
- Cosmic ménage à trois: the origin of satellite galaxies on extreme orbits. MNRAS 379 (4), pp. 1475–1483. External Links: Document, 0704.1773, ADS entry Cited by: §3.3.3.
- Measuring bulk flows of the intracluster medium in the Perseus and Coma galaxy clusters using XMM-Newton. A&A 633, pp. A42. External Links: Document, 1911.13108, ADS entry Cited by: §1, §3.1.
- XMM-Newton Observation of IC 310 in the Outer Region of the Perseus Cluster of Galaxies. PASJ 57, pp. 743–749. External Links: Document, astro-ph/0509386, ADS entry Cited by: §3.1.
- The structure of the intracluster medium of the Perseus cluster.. A&A 256, pp. L11–L14. External Links: ADS entry Cited by: §1, §2.4, §3.1, §3.1.
- A New Class of X-Ray Tails of Early-type Galaxies and Subclusters in Galaxy Clusters: Slingshot Tails versus Ram Pressure Stripped Tails. ApJ 874 (2), pp. 112. External Links: Document, 1903.00482, ADS entry Cited by: §3.3.1.
- Multifrequency radio continuum observations of head-tail galaxies in the Perseus cluster. A&A 331, pp. 901–915. External Links: ADS entry Cited by: §3.3.1, §3.3.2, §3.3.2.
- Large-scale Motions in the Perseus Galaxy Cluster. ApJ 757 (2), pp. 182. External Links: Document, 1208.2990, ADS entry Cited by: §1, §2.4, §3.1, §3.1, §3.1.
- Baryons at the Edge of the X-ray-Brightest Galaxy Cluster. Science 331 (6024), pp. 1576. External Links: Document, 1102.2429, ADS entry Cited by: §1.
- The cosmological simulation code GADGET-2. MNRAS 364 (4), pp. 1105–1134. External Links: Document, astro-ph/0505010, ADS entry Cited by: §3.3.4.
- SRG X-ray orbital observatory. Its telescopes and first scientific results. A&A 656, pp. A132. External Links: Document, 2104.13267, ADS entry Cited by: §1, §2.
- Gas Bulk Motion in the Perseus Cluster Measured with Suzaku. ApJ 782 (1), pp. 38. External Links: Document, 1312.4526, ADS entry Cited by: §1, §2.4.
- LOFAR high-band antenna observations of the Perseus cluster: The discovery of a giant radio halo. A&A 692, pp. A12. External Links: Document, 2410.02863, ADS entry Cited by: Figure 7, §3.3.1, §3.3.2, §3.3.
- Chandra Cluster Cosmology Project. II. Samples and X-Ray Data Reduction. ApJ 692 (2), pp. 1033–1059. External Links: Document, 0805.2207, ADS entry Cited by: Appendix A.
- Is There an Enormous Cold Front at the Virial Radius of the Perseus Cluster?. ApJ 929 (1), pp. 37. External Links: Document, 2006.14043, ADS entry Cited by: §1, §2.4, §3.2.
- Disentangling Multiple Gas Kinematic Drivers in the Perseus Galaxy Cluster. arXiv e-prints, pp. arXiv:2509.04421 (Accepted for publication in Nature). External Links: Document, 2509.04421, ADS entry Cited by: §4.
- The Interaction of Matter and Radiation in a Hot-Model Universe. Ap&SS 4 (3), pp. 301–316. External Links: Document, ADS entry Cited by: §2.1.
- Collision of merger and accretion shocks: formation of Mpc-scale contact discontinuity in the Perseus cluster. MNRAS 498 (1), pp. L130–L134. External Links: Document, 2007.02551, ADS entry Cited by: §1, §3.2.
- Runaway merger shocks in galaxy cluster outskirts and radio relics. MNRAS 488 (4), pp. 5259–5266. External Links: Document, 1904.03052, ADS entry Cited by: §3.3.1.
- Mapping the Perseus Galaxy Cluster with XRISM: Gas Kinematic Features and their Implications for Turbulence. arXiv e-prints, pp. arXiv:2510.12782 (Accepted for publication in A&A). External Links: Document, 2510.12782, ADS entry Cited by: §4.
- A shock near the virial radius of the Perseus Cluster. A&A 652, pp. A147. External Links: Document, 2107.01663, ADS entry Cited by: §1, §2.4, §3.2.
- Gas density fluctuations in the Perseus Cluster: clumping factor and velocity power spectrum. MNRAS 450 (4), pp. 4184–4197. External Links: Document, 1501.07271, ADS entry Cited by: §1.
- Cold fronts: probes of plasma astrophysics in galaxy clusters. Journal of Plasma Physics 82 (3), pp. 535820301. External Links: Document, 1603.08882, ADS entry Cited by: §1.
Appendix A Chandra, XMM-Newton, and SRG/eROSITA datasets